# Geodesic Slice Sampling on the Sphere

### v3

@@ Appendix D Proof of Lemma 11
×$\quad\hskip 170.71652pt\times\bar{Q}_{L(T_{\theta}(x,v),t)}(0,\gamma_{T_{\theta}(x,v)}^{-1}(A))\ \mu_{x}({\rm d}v)\,\sigma_{d-1}({\rm d}x)\,{\rm d}\theta.$
@@
$=\int_{L(t)}\int_{{\mathbb{S}^{d-1}_{x}}}\frac{1}{{\lambda}(L(x,v,t))}\int_{\gamma_{(x,v)}^{-1}(B)\cap L(x,v,t)}\bar{Q}_{L(x,v,t)}(\theta,\gamma_{(x,v)}^{-1}(A)){\rm d}\theta\,\mu_{x}({\rm d}v)\,\sigma_{d-1}({\rm d}x).$
@@
$\qquad=\int_{L(t)}\int_{{\mathbb{S}^{d-1}_{x}}}\frac{1}{{\lambda}(L(x,v,t))}\int_{\gamma_{(x,v)}^{-1}(A)\cap L(x,v,t)}\bar{Q}_{L(x,v,t)}(\theta,\gamma_{(x,v)}^{-1}(B)){\rm d}\theta\,\mu_{x}({\rm d}v)\,\sigma_{d-1}({\rm d}x).$
@@
Since RWMH suffers from diffusive behavior, we also test spherical Hamiltonian Monte Carlo (HMC) suggested by Lan et al. (2014) as an alternative MCMC approach. In spherical HMC, the sample space is first augmented by momenta or velocities $v\in\mathbb{R}^{d}$ that are in the the tangent space $E_{x}$ to $\mathbb{S}^{d-1}$ at the current sample $x$. The momenta follow a standard Normal distribution, and the Markov chain is generated in the space $\{(x,v)\mid x\in\mathbb{S}^{d-1},v\in E_{x}\}$. Diffusive behavior is suppressed by the special type of proposal step that solves Hamilton’s equations of motion by using a leapfrog integrator (Neal, 2011). During leapfrog integration, the gradient of the log probability $\nabla\log p(x)$ guides the Markov chain, thereby reaching nearby modes in much shorter time than RWMH. The spherical HMC algorithm is detailed in Algorithm 8. In contrast to standard HMC, spherical HMC involves a rotation of the positions and momenta during leapfrog integration where the rotation matrix is a Givens rotation $G(v/\|v\|,x,\theta)$ in the plane spanned by the momenta $v$ and the positions $x$ (see lines 8–10 in Algorithm 8 and Equation 4). In addition to the step-size parameter $\varepsilon>0$, we also need to choose the number of leapfrog steps $T\in\mathbb{N}$. In our experiments, we always set $T=10$.
@@ Appendix H Step-size Tuning for RWMH and HMC
Both RWMH and spherical HMC involve a step-size parameter $\varepsilon$. Because a good choice of $\varepsilon$ depends on the particular shape of our target distribution $p(x)$, we first find $\varepsilon$ automatically during a burn-in phase. During burn-in, we increase $\varepsilon$ by a factor 1.02, if the proposal (based on the current value of $\varepsilon$) is accepted. We decrease $\varepsilon$ by a factor 0.98 if the proposed state is rejected. After a burn-in phase, the value of $\varepsilon$ is kept fixed.
Algorithm 8 Hamiltonian Monte Carlo on $\mathbb{S}^{d-1}$ for step-size $\varepsilon>0$ and number of integration steps $T$
@@
Appendix I Tests on 50-dimensional Bingham Distribution.
Tests for the 50-dimensional Bingham distribution with a concentration parameter $\kappa_{50}=300$.
Figure A1: Histograms of approximate Bingham samples ($d=50,\kappa_{d}=300$) projected on the first mode obtained with each MCMC method are shown in grey. The red dashed line indicates the baseline obtained with the acceptance/rejection sampler of Kent et al. (2018).
Figure A2: Left: Autocorrelation analysis of $x\mapsto u_{d}^{T}x$ with respect to approximate samples from a Bingham distribution with $d=50$ and $\kappa_{d}=300$. Right: Estimated hopping frequency between both modes of the Bingham distribution. Note that the hopping frequency is shown on a logarithmic axis.
Figure A3: Geodesic distance (log-scale) between successive approximate samples for a Bingham distribution with $d=50$ and $\kappa_{d}=300$.
Appendix J Run Times of Numerical Illustrations
We summarize the run times in seconds (s) for all the numerical experiments tested on Intel i7-10510U CPU in Table A1. We nomenclate these experiments for brevity. For the Bingham distribution with $d=10$, $\kappa_{d}=30$ and $N=10^{5}$ as “Bingham-d10”, for the Bingham distribution with $d=50$, $\kappa_{d}=300$ and $N=10^{5}$ as “Bingham-d50”, for the mixture of vMF distribution with $d=10$, $\kappa=100$, $K=5$ and $N=10^{6}$ as “vMF-d10” and for the curve vMF distribution with $\kappa=300$ and $N=10^{5}$ as “curved-vMF”.

### v4

@@ Appendix D Proof of Lemma 11
$\quad\hskip 170.71652pt\times\bar{Q}_{L(T_{\theta}(x,v),t)}(0,\gamma_{T_{\theta}(x,v)}^{-1}(A))\ \mu_{x}({\rm d}v)\,\sigma_{d-1}({\rm d}x)\,{\rm d}\theta.$
@@
$=\int_{L(t)}\int_{{\mathbb{S}^{d-2}_{x}}}\frac{1}{{\lambda}(L(x,v,t))}\int_{\gamma_{(x,v)}^{-1}(B)\cap L(x,v,t)}\bar{Q}_{L(x,v,t)}(\theta,\gamma_{(x,v)}^{-1}(A)){\rm d}\theta\,\mu_{x}({\rm d}v)\,\sigma_{d-1}({\rm d}x).$
@@
$\qquad=\int_{L(t)}\int_{{\mathbb{S}^{d-2}_{x}}}\frac{1}{{\lambda}(L(x,v,t))}\int_{\gamma_{(x,v)}^{-1}(A)\cap L(x,v,t)}\bar{Q}_{L(x,v,t)}(\theta,\gamma_{(x,v)}^{-1}(B)){\rm d}\theta\,\mu_{x}({\rm d}v)\,\sigma_{d-1}({\rm d}x).$
@@
Since RWMH suffers from diffusive behavior, we also test spherical Hamiltonian Monte Carlo (HMC) suggested by Lan et al. (2014) as an alternative MCMC approach. In spherical HMC, the sample space is first augmented by momenta or velocities $v\in\mathbb{R}^{d}$ that are in the the tangent space $E_{x}$ to $\mathbb{S}^{d-1}$ at the current sample $x$. The momenta follow a standard Normal distribution, and the Markov chain is generated in the space $\{(x,v)\mid x\in\mathbb{S}^{d-1},v\in E_{x}\}$. Diffusive behavior is suppressed by the special type of proposal step that solves Hamilton’s equations of motion by using a leapfrog integrator (Neal, 2011). During leapfrog integration, the gradient of the log probability $\nabla\log p(x)$ guides the Markov chain, thereby reaching nearby modes in much shorter time than RWMH. The spherical HMC algorithm is detailed in Algorithm 8. In contrast to standard HMC, spherical HMC involves a rotation of the positions and momenta during leapfrog integration where the rotation matrix is a Givens rotation $G(v/\|v\|,x,\theta)$ in the plane spanned by the momenta $v$ and the positions $x$ (see lines 8–10 in Algorithm 8 and (4)). In addition to the step-size parameter $\varepsilon>0$, we also need to choose the number of leapfrog steps $T\in\mathbb{N}$. In our experiments, we always set $T=10$.
@@ Appendix H Step-size Tuning for RWMH and HMC
Both RWMH and spherical HMC involve a step-size parameter $\varepsilon$. Because a good choice of $\varepsilon$ depends on the particular shape of our target distribution, we first find $\varepsilon$ automatically during a burn-in phase. During burn-in, we increase $\varepsilon$ by a factor 1.02, if the proposal (based on the current value of $\varepsilon$) is accepted. We decrease $\varepsilon$ by a factor 0.98 if the proposed state is rejected. After a burn-in phase, the value of $\varepsilon$ is kept fixed.
Algorithm 8 Hamiltonian Monte Carlo on $\mathbb{S}^{d-1}$ for step-size $\varepsilon>0$ and number of integration steps $T$ .
@@
Appendix I Run Times of Numerical Illustrations
We summarize the run times in seconds (s) for all the numerical experiments tested on Intel i7-10510U CPU in Table A1. In addition to the tests reported in Section 4, we provide a Jupyter notebook for testing the geodesic slice samplers on Bingham targets in various dimensions.111github.com/microscopic-image-analysis/geosss/scripts/Bingham.ipynb We also report run times for these targets. For brevity we name these experiments “Bingham-d10” for the Bingham distribution with $d=10$, $\kappa_{d}=30$ and $N=10^{5}$; “Bingham-d50” for the Bingham distribution with $d=50$, $\kappa_{d}=300$ and $N=10^{5}$; “vMF-d10” for the mixture of vMF distribution with $d=10$, $\kappa=100$, $K=5$ and $N=10^{6}$; “curved-vMF-d3” for the curve vMF distribution with $d=3$, $\kappa=500$ and $N=10^{6}$; “curved-vMF-d5” for curve vMF with $d=5$, $\kappa=800$ and $N=10^{6}$.